OpenStudy (anonymous):

what is the difference in simplest form? n^2+3n+2/n^2+6b+8 - 2n/n+4

OpenStudy (anonymous):

\[\frac{ n^2+3n+2 }{ (n+4)(n+2) }-\frac{ 2n }{ n+4 }=\frac{ n^2+3n+2 }{ (n+4)(n+2) }-\frac{ 2n(n+2) }{(n+4)(n+2) }\]\[=\frac{ n^2+3n+2 }{ (n+4)(n+2) }-\frac{ 2n^2+4n }{(n+4)(n+2) }=\frac{ (n^2+3n+2)-(2n^2+4n) }{ (n+4)(n+2) }\]\[=\frac{ -n^2-n+2}{ (n+4)(n+2) }=\frac{ -(n+2)(n-1)}{ (n+4)(n+2) }\]\[=-\frac{ n-1}{ n+4 }\] @OSMiUM_MONkEY

OpenStudy (anonymous):

1-n?

OpenStudy (anonymous):

No, after you subtract, the final answer is:\[-\frac{ n-1 }{ n+4 }\] @OSMiUM_MONkEY

OpenStudy (anonymous):

Get it? And btw, don't forget the medal =D

OpenStudy (anonymous):

@OSMiUM_MONkEY